. A flooring tile has the shape of a parallelogram whose base is 24 cm and the corresponding height is 10 cm. How many such tiles are required to cover a floor of area 1080 m^2? [If required you can split the tiles in whatever way you want to fill up the corners]

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given: Base of flooring tile = 24 cm = 0.24 m
Corresponding height of a flooring tile= 10 cm = 0.10 m
Now Area of flooring tile= Base×Altitude
= 0.24×0.10
= 0.024
Area of flooring tile is 0.024m^2
Number of tiles required to cover the floor= Area of floor/Area of one tile = 1080/0.024 = 45000 tiles

Thanks